# Web Archives Metadata Generation with gpt-4o: Challenges and Insights

### v1

@@ 1 Introduction
The digital landscape is constantly evolving and the need to preserve our online heritage has become increasingly urgent. The Resource Discovery (RD) department of the National Library Board Singapore (NLB) [1] provides cataloguing services for the collections of the National Library (NL) and public libraries in Singapore. NL, like many other archives and libraries worldwide, collects and archives websites in its effort to preserve the mercurial history of the web. This collection is known as “Web Archive Singapore” [2]. Each website is manually reviewed and catalogued according to the local application profile based on Dublin Core [3]. NL had been crawling websites on a curated basis since 2006, with each website requiring individual website owner consent. In 2019, the National Library Board Act [4] was amended to impart NL with the authority to harvest all websites in the .sg domain without explicit consent from each individual website owner. The legislative change resulted in the explosive growth of the web collection [5]. As of June 2024, there are over 98,000 unique websites and webpages from both the domain crawl and curated collecting [2]. With the large collection numbers, RD sought to explore if technology such as generative AI would be viable to facilitate the cataloguing of individual sites harvested in the yearly domain crawl. RD wished to explore the generation of titles, abstracts and subjects for each website as these metadata fields were deemed the most crucial and used for display and discovery.
@@ 1.2 Problem Statement
    Efficiency: How can we develop an automated system that can process and generate metadata for a large number of websites that significantly reduces the time and resources required?
@@
These challenges are non-trivial for collections such as WAS, which aims to preserve the digital heritage of Singaporean life, culture, and history in the 21st century [2]. With an estimated 20,000 Web ARChive (WARC) files created annually that comply with quality control standards, there is a critical need for a scalable, reliable, and cost-effective solution for metadata generation.
@@ 2.2 Metadata Strategies for Web Archives
This work focuses on the provision of descriptive metadata to support search and access of individual sites in a web archive. There can be varying metadata approaches [11] such as :
  • Harvesting the metadata as-is
  • Reviewing only basic metadata fields such as Title and Language
  • Providing fuller metadata for specially curated web collections, including assigning subject headings
In many cases, a combination of approaches is used. For institutions conducting large scale automating crawling (e.g. by domain), this would be:
  • for websites that have been curated into themes, metadata is individually reviewed and enhanced
  • for the remainder of full crawl, metadata is used as-is
@@
Cataloguing every single website in the domain crawl is a time-consuming task and given the explosive progress and availability of AI/ML tools in recent years, especially with Large Language Models (LLMs), we were keen to explore the possibilities of integrating the technology into our workflow.
Libraries have naturally explored myriad use cases for AI/ML. In the case of cataloguing, there have been many attempts at using generative AI for cataloguing and metadata enhancement [13, 14].The IFLA WLIC 2023 programme featured an entire session on “Utopia, threat or opportunity first? Artificial Intelligence and machine learning for cataloguing”, with speakers from four different countries sharing on the use of AI/ML in cataloguing applications, such as generating catalogue records, detecting metadata from scanned resources, and deduplication [15].
While results have been promising, most projects embody what was said by Allen [16] in a Library of Congress blog post (emphasis ours):
“With all the possibilities, why aren’t we already embedding machine learning into everything we do? The answer is complicated, but at its simplest, machine learning (ML) and artificial intelligence (AI) tools haven’t demonstrated that they’re able to meet our very high standards for responsible stewardship of information in most cases, without significant human intervention”.
Interest in AI/ML has naturally extended to the web archive community. This is embodied in the program for the 2024 IIPC Web Archiving Conference, which featured a dedicated session themed “AI & Machine Learning”, in addition to having other Lightning and Drop-In talks that featured AI/ML use  [17]. We know there are many projects utilising AI/ML for web archives such as the Harvard Law School Library Innovation Lab’s WARC-GPT, launched in February 2024, which is a custom chatbot for users to explore a WARC collections via natural language queries  [18]. To date, however, we have not come across specific work on the application of AI/ML for cataloguing by directly processing WARC files.
@@ 2.3 Large Language Models
The advent of LLMs has revolutionized the field of Natural Language Processing (NLP). These sophisticated models, capable of processing and generating human-like text, have surpassed traditional methods in numerous NLP tasks. What was once considered a complex challenge, such as text summarization, is now routinely managed with impressive accuracy and efficiency by LLMs. Leading the charge in the commercialization of large language models are models such as gpt-4o by OpenAI [19], Claude by Anthropic [20], and Gemini by Google [21]. Built upon the groundbreaking transformer architecture introduced in "Attention Is All You Need" [22], these models leverage the attention mechanism to outperform earlier methods for NLP tasks that relied heavily on linguistic techniques like stemming and lemmatization or other neural networks like Recurrent Neural Networks (RNN) and Gated Recurrent Units (GRU). The integration of Reinforcement Learning from Human Feedback [23] and Direct Preference Optimization [24] further enhances the quality of LLM-generated output, making them more human-like thus suitable for generation of high quality content. Fine-tuning LLMs with datasets from a particular domain not only boosts their contextual understanding and expertise in domain specific tasks but also enhances the accuracy of its output [25, 26].
@@ 2.4 In-Context Learning
However, given the limitations of time and computational resources, the process of fine-tuning Language Models for specific downstream tasks often incurs significant costs. As a result of this challenge, in-context learning (ICL) has gained significant popularity. The ICL technique enables LLMs to acquire new tasks without modifying the model parameters, by relying on task-specific examples provided in the input context. ICL has demonstrated strong performance in few-shot and zero-shot learning. Zero-shot learning leverages the model’s existing knowledge to deduce task requirements. As illustrated by Brown et al., GPT-3 was utilized to showcase the system’s versatility in the absence of task-specific fine-tuning. With few-shot learning, the model is provided with a restricted number of examples to showcase the task, resulting in enhanced comprehension and performance [8]. Radford et al have demonstrated that few-shot prompting can significantly enhance the model’s accuracy across a diverse range of tasks [27].
@@
Multi-query reasoning techniques, like Graph of Thought (GoT) [29], Tree of Thought (ToT) [30], and Least-to-Most [31], use multiple LLM queries to extract various tenable reasoning paths. By breaking down complicated issues into smaller, more manageable subproblems, these techniques improve the reasoning power of the model.In conclusion, in-context learning strategies are essential for fully utilizing LLMs. The rationale behind us selecting CoT is its capacity to minimize token consumption while yielding superior outcomes.
@@ 3.1 Data Collection and Preparation
To obtain web data for our inquiry, we collected a total of 112 WARC (Web ARChive) files from the Web Archives of Singapore. The HTML content from these files was extracted using the WARCIO [32] and fastwarc [33] python libraries. BeautifulSoup facilitated the extraction of relevant metadata, such as titles and primary text content, from the HTML. We removed unnecessary tags and scripts during this procedure to ensure that the main content is highlighted. We standardized the URLs and conducted a quality assurance assessment to eliminate any substandard or irrelevant data, ensuring uniformity. This involved identifying common indicators of non-functional material, such as "404" errors or placeholder text like "lorem ipsum." In addition, we implemented a deduplication technique to consolidate individual records obtained from duplicate URLs. This ensured the preservation of the information’s originality and relevance.
We employed multi-threaded processing using a Thread PoolExecutor to efficiently handle the vast amount of data, resulting in optimized resource utilization and significantly reduced processing time. The final step was combining the processed data into DataFrames, which were then ready for further analysis and model development. An indispensable requirement for the reliability and robustness of our investigation was a high-caliber dataset ensured by this approach.
@@ 3.2 Heuristics for Data Reduction
To reduce the number of tokens (and consequently cost) of the input to be used, our methodology involved developing and evaluating various heuristic methods for efficient content extraction and summarization. The three heuristics were:
@@
    About Page Priority: We prioritized extracting content from the ’About’ page of the website. In cases where an ’About’ page was unavailable, we defaulted to using the content from the shortest URL
@@
    Shortest URL: This method involved extracting content exclusively from the web page with the shortest URL
@@ 3.3 Prompt Engineering for Title and Abstract Generation
Prompt 2:
Use the previous prompt and add:
@@ 3.4 Evaluation Methods
  • Levenshtein Distance [34] for title comparison
  • BERTScore [35](using bert-based-cased as the embedding) for title and content similarity.
@@
Our ranking algorithm evaluates heuristic variants based on their ability to generate abstracts and titles through an aggregated scoring system. This method integrates three primary criteria: the minimal median of Levenshtein Distance, the maximum median of BERTScore, and the minimum standard deviation. The semantic similarity between reference and produced texts is assessed by BERTScore through contextual embeddings from BERT; a larger median indicates increased quality. The exact match accuracy is measured by the Levenshtein Distance; a lower median indicates a closer alignment with the reference. The standard deviation of BERTScores is a metric that indicates consistency; values that are lower indicate more consistent performance. A total score is generated by combining the rankings of heuristics for each criterion. This comprehensive method ensures a fair evaluation of accuracy and consistency, thereby influencing our selection of heuristics for text production. The 2D and 3D visualizations of BERTScore of our generated data is demonstrated in Figure 1 and Figure 3. In the 2D BERT embedding vector space, it is evident that certain abstract pairs and title pairs are identical. The majority of the distances between abstract pairs is negligible.
Figure 3: Bar chart of ranked aggregated scores for the 6 different combinations of prompts and heuristics. The scores are based on criteria including maximum BERTScore median, minimum Levenshtein median, and minimum standard deviation.
@@
Table 1: A table used for comparison of ranked aggregated scores for heuristic and prompt combinations with and without Rules
@@
Manual Evaluation: Based on the automated evaluation results, we shortlisted the two approaches for manual evaluation. These titles and abstracts, in combination with a set generated by a human, were then subjected to manual grading by a team of eight trained cataloguers. The results were evaluated using the statistical tests. Cochran’s Q Test [36] and McNemar’s Test [37] are statistical tools that are employed to assess categorical data, particularly in situations involving related samples or repeated measurements. These tests are employed to compare the performance of human evaluation and various heuristics in our research.
McNemar’s Test is often used to analyze nominal data that is paired, particularly in before-and-after studies or when comparing two related samples. It is particularly beneficial for binary results in matched pairs of subjects. In this study, we employ McNemar’s Test to evaluate heuristic 2 with human assessment. This pairwise comparison enables us to determine whether the automated heuristic and human judgment exhibit statistically significant differences.
Cochran’s Q Test is an expanded version of McNemar’s test that is applicable to more than two related groups. It is implemented when three or more conditions that are related generate binary outcomes, such as "pass" or "fail." The test assesses whether the proportions of these binary outcomes vary statistically significantly across the conditions. We utilize Cochran’s Q Test to determine whether the pass/fail rates of the three evaluation approaches(heuristic 2, heuristic 3, and human) display any significant variations. This test elucidates whether the approaches are consistently producing distinct results or if the variations we observe could be the result of random chance.
In our dataset, there are numerous evaluations of the same set of objects, including both human and heuristic evaluations. These tests are intended for samples of this nature that are related. Both tests are qualified by our pass/fail (1/0) data. The purpose of these tests is to conduct a precise statistical comparison of the performance of various evaluation techniques. The McNemar’s Test enables us to focus on the comparison between heuristic 2 and human evaluation, which may be particularly helpful in assessing the automated approach’s performance against human judgment.
@@ 3.5 Software and Hardware
To verify if heuristics are able to effectively reduce the number of tokens needed for metadata generation, we wrote a python package warc2summary which provides a pipeline for metadata creation and evaluation.In short, this package uses WARCIO [32] and FastWARC [33] as the WARC file processor and for title and abstract generation we use OpenAI API (gpt-4o) as the main model and Instructor [38] to constrain the output.
We ran the software on a Lenovo Thinkpad T14 with a 11th Gen Intel(R) Core(TM) i5-1145G7 CPU and 16GB of DDR4 RAM. The pipeline took close to 2 hours to process the WARC files and another 2 hours to generate the synthetic data from API calls. We expect faster performance in WARC file processing with a better CPU. In addition, the API call pipelines were called sequentially in order to not breach OpenAI API rate limits. With higher rate limits, it is possible to parallelise the calls for better performance.The WARC files are warc.gz files and are taken from Web Archive Singapore. This provided us with a Dataframe containing synthetic titles and abstracts.
We identified 6 different combinations of prompts and heuristics as stated above. We then used a ranked aggregation scoring method, incorporating Levenshtein Distance for the title and BERTScore (using bert-based-cased embeddings), with criteria of maximum BERTScore median, minimum Levenshtein median, and minimum standard deviation for both, to shortlist two specific combinations of prompts and heuristics. Finally, a team of 8 trained cataloguers performed manual grading.

### v3

@@ 1 Introduction
The digital landscape is constantly evolving, and the need to preserve our online heritage has become increasingly urgent. The Resource Discovery (RD) department of the National Library Board Singapore (NLB) provides cataloguing services for the collections of the National Library (NL) and public libraries in Singapore. NL, like many other archives and libraries worldwide, collects and archives websites in its effort to preserve the mercurial history of the web [1]. This collection is known as Web Archive Singapore [2]. Each website is manually reviewed and catalogued according to the local application profile based on Dublin Core [3]. NL has been crawling websites on a curated basis since 2006, with each website requiring individual website owner consent. In 2019, the National Library Board Act was amended to impart NL with the authority to harvest all websites in the .sg domain without explicit consent from each individual website owner [4]. The legislative change resulted in the explosive growth of the web collection [5]. As of June 2024, there are over 98,000 unique websites and webpages from both the domain crawl and curated collection [2]. With the large collection numbers, RD sought to explore whether technology such as generative AI would be viable to facilitate the cataloguing of individual sites harvested in the yearly domain crawl. RD wished to explore the generation of titles, abstracts, and subjects for each website, as these metadata fields were deemed the most crucial and used for display and discovery.
@@ 1.2 Problem Statement
    Efficiency: How can we develop an automated system that can process and generate metadata for a large number of websites, significantly reducing the time and resources required?
@@
These challenges are nontrivial for collections such as WAS, which aims to preserve the digital heritage of Singaporean life, culture, and history in the 21st century. With an estimated 20,000 Web ARChive (WARC) files created annually that comply with quality control standards, there is a critical need for a scalable, reliable, and cost-effective solution for metadata generation.
@@ 2.1 Metadata Strategies for Web Archives
This work focuses on the provision of descriptive metadata to support search and access of individual sites in a web archive. There can be varying metadata approaches such as  [11]:
  • harvesting the metadata as-is;
  • reviewing only basic metadata fields such as ’Title’ and ’Language’, or
  • providing fuller metadata for specially curated web collections, including assigning subject headings [11]
In many cases, a combination of approaches is used. For institutions conducting large-scale automated crawling (e.g., by domain), this would be:
  • for websites that have been curated into themes, metadata is individually reviewed and enhanced;
  • for the remainder of the full crawl, metadata is used as-is;
@@
Cataloguing every single website in the domain crawl is a time-consuming task and given the explosive progress and availability of artificial intelligence/machine learning (AI/ML) tools in recent years, especially with large language models (LLMs), we were keen to explore integrating this technology into our workflow.
Libraries have naturally explored myriad use cases for AI/ML. In the case of cataloguing, there have been many attempts at using generative AI for cataloguing and metadata enhancement [13, 14].The IFLA WLIC 2023 programme featured an entire session on “Utopia, Threat or Opportunity First? Artificial Intelligence and Machine Learning for Cataloguing,” with speakers from four different countries sharing on the use of AI/ML in cataloguing applications, such as generating catalogue records, detecting metadata from scanned resources, and deduplication [15].
While results have been promising, most projects embody what was said by Allen in a Library of Congress blog post:
"With all the possibilities, why aren’t we already embedding machine learning into everything we do? The answer is complicated, but at its simplest, machine learning (ML) and artificial intelligence (AI) tools haven’t demonstrated that they’re able to meet our very high standards for responsible stewardship of information in most cases, without significant human intervention"[16].
Interest in AI/ML has naturally extended to the web archive community. This was embodied in the program for the 2024 IIPC Web Archiving Conference, which featured a dedicated session themed “AI & Machine Learning,” in addition to having other lightning and drop-in talks that featured AI/ML use [17]. We know of many projects utilizing AI/ML for web archives such as the Harvard Law School Library Innovation Lab’s WARC-GPT, launched in February 2024, which is a custom chatbot for users to explore WARC collections via natural language queries [18]. To date, however, we have not come across specific work on the application of AI/ML for cataloguing by directly processing WARC files.
@@ 2.2 Large Language Models
The advent of LLMs has revolutionized the field of natural language processing (NLP). These sophisticated models, capable of processing and generating human-like text, have surpassed traditional methods in numerous NLP tasks. What was once considered a complex challenge, such as text summarization, is now routinely managed with impressive accuracy and efficiency by LLMs. Leading the charge in the commercialization of large language models are models such as GPT-4o by OpenAI, Claude by Anthropic, and Gemini by Google [19, 20, 21]. Built upon the groundbreaking transformer architecture introduced in “Attention Is All You Need,” these models leverage the attention mechanism to outperform earlier methods for NLP tasks that relied heavily on linguistic techniques like stemming and lemmatization or other neural networks like recurrent neural networks (RNN) and gated recurrent units (GRU) [22]. The integration of reinforcement learning from human feedback and direct preference optimization further enhances the quality of LLM-generated output, making them more human-like and thus suitable for generation of high-quality content [23, 24]. Fine-tuning LLMs with datasets from a particular domain not only boosts their contextual understanding and expertise in domain specific tasks but also enhances the accuracy of their output [25, 26].
@@ 2.3 In-Context Learning
However, given the limitations of time and computational resources, the process of fine-tuning language models for specific downstream tasks often incurs significant costs. As a result of this challenge, in-context learning (ICL) has gained significant popularity. The ICL technique enables LLMs to acquire new tasks without modifying the model parameters, by relying on task-specific examples provided in the input context. ICL has demonstrated strong performance in few-shot and zero-shot learning. Zero-shot learning leverages the model’s existing knowledge to deduce task requirements. As illustrated by Brown et al., GPT-3 was utilized to showcase the system’s versatility in the absence of task-specific fine-tuning. With few-shot learning, the model is provided with a restricted number of examples to showcase the task, resulting in enhanced comprehension and performance. Radford et al. have demonstrated that few-shot prompting can significantly enhance the model’s accuracy across a diverse range of tasks[27].
@@
Multi-query reasoning techniques, like graph of thought (GoT), tree of thought (ToT), and least-to most, use multiple LLM queries to extract various tenable reasoning paths [29, 30, 31]. By breaking down complicated issues into smaller, more manageable subproblems, these techniques improve the reasoning power of the model. In conclusion, in-context learning strategies are essential for fully utilizing LLMs. The rationale behind us selecting CoT is its capacity to minimize token consumption while yielding superior outcomes.
@@ 3.1 Data Collection and Preparation
To obtain web data for our inquiry, we collected a total of 112 WARC (Web ARChive) files from the Web Archives of Singapore. The HTML content from these files was extracted using the WARCIO and fastwarc python libraries [32, 33]. BeautifulSoup facilitated the extraction of relevant metadata, such as titles and primary text content, from the HTML. We removed unnecessary tags and scripts during this procedure to ensure that the main content is highlighted. We standardized the URLs and conducted a quality assurance assessment to eliminate any substandard or irrelevant data, ensuring uniformity. This involved identifying common indicators of nonfunctional material, such as “404” errors or placeholder text like “lorem ipsum.” In addition, we implemented a deduplication technique to consolidate individual records obtained from duplicate URLs. This ensured the preservation of the information’s originality and relevance. We employed multithreaded processing using a ThreadPoolExecutor to efficiently handle the vast amount of data, resulting in optimized resource utilization and significantly reduced processing time. The final step was combining the processed data into DataFrames, which were then ready for further analysis and model development. An indispensable requirement for the reliability and robustness of our investigation was a high caliber dataset ensured by this approach. The final step was combining the processed data into DataFrames, which were then ready for further analysis and model development. An indispensable requirement for the reliability and robustness of our investigation was a high-caliber dataset ensured by this approach.
@@ 3.2 Heuristics for Data Reduction
To reduce the number of input tokens and thereby achieve cost savings, our methodology involved developing and evaluating various heuristic methods for efficient content extraction and summarization. The three heuristics were:
@@
    About Page Priority: We prioritized extracting content from the “About” page. If no “About” page existed, we used content from the shortest URL.
@@
    Shortest URL: We extracted content exclusively from the web page with the shortest URL.
@@ 3.3 Prompt Engineering for Title and Abstract Generation
Prompt 2 (appended to Prompt 1):
@@ 3.4 Evaluation Methods
  • Levenshtein distance for title comparison  [34]
  • BERTScore (using bert-based-cased as the embedding) for title and content similarity  [35]
@@
Our ranking algorithm evaluates heuristic variants based on their ability to generate abstracts and titles through an aggregated scoring system. This method integrates three primary criteria: the minimal median of Levenshtein distance, the maximum median of BERTScore, and the minimum standard deviation. The semantic similarity between reference and produced texts is assessed by BERTScore through contextual embeddings from BERT; a larger median indicates increased quality. The exact match accuracy is measured by the Levenshtein distance; a lower median indicates a closer alignment with the reference. The standard deviation of BERTScores is a metric that indicates consistency; values that are lower indicate more consistent performance. A total score is generated by combining the rankings of heuristics for each criterion. This comprehensive method ensures a fair evaluation of accuracy and consistency, thereby influencing our selection of heuristics for text production. The 2D and 3D visualizations of BERTScore of our generated data are demonstrated in figure 1 and figure 3. In the 2D BERT embedding vector space, it is evident that certain abstract pairs and title pairs are identical. The majority of the distances between abstract pairs is negligible.
Figure 3: Bar chart of ranked aggregated scores for the six different combinations of prompts and heuristics. The scores are based on criteria including maximum BERTScore median, minimum Levenshtein median, and minimum standard deviation.
@@
Table 1: Comparison of ranked aggregated scores for different prompt and heuristic combinations. The highlighted rows, heuristic 2 with rules and heuristic 3 without rules, achieved the best overall performance, each with a top-ranked score of 1. These combinations were selected for manual evaluation based on their superior results in automated scoring metrics.
@@
Manual evaluation:
Based on the automated evaluation results, we shortlisted the two approaches for manual evaluation. These titles and abstracts, in combination with a set generated by a human, were then subjected to manual grading by a team of eight trained cataloguers. The results were evaluated using the statistical tests. Cochran’s Q test and McNemar’s test are statistical tools that are employed to assess categorical data, particularly in situations involving related samples or repeated measurements [36, 37]. These tests are employed to compare the performance of human evaluation and various heuristics in our research.
McNemar’s test is often used to analyze nominal data that is paired, particularly in before-and-after studies or when comparing two related samples. It is particularly beneficial for binary results in matched pairs of subjects. In this study, we employ McNemar’s test to evaluate combination 2 with human assessment. This pairwise comparison enables us to determine whether the automated heuristic and human judgment exhibit statistically significant differences. Cochran’s Q test is an expanded version of McNemar’s test that is applicable to more than two related groups. It is implemented when three or more conditions that are related generate binary outcomes, such as “pass” or “fail.” The test assesses whether the proportions of these binary outcomes vary statistically significantly across the conditions. We apply Cochran’s Q test to determine whether the pass/fail rates of the three evaluation approaches (combination 2, combination 3, and human) display any significant variations. This test elucidates whether the approaches are consistently producing distinct results or if the variations could be the result of random chance.
In our dataset, there are numerous evaluations of the same set of objects, including both human and heuristic evaluations. Both tests are qualified by our pass/failure data. The purpose of these tests is to conduct a precise statistical comparison of the performance of various evaluation techniques. McNemar’s test enables us to focus on the comparison between two combinations and human evaluation, which may be helpful in assessing the automated approach’s performance against human judgment.
@@ 3.5 Software and Hardware
To verify if heuristics can effectively reduce the number of tokens needed for metadata generation, we wrote a python package warc2summary which provides a pipeline for metadata creation and evaluation. In short, this package uses WARCIO and fastwarc as the WARC file processor, and for title and abstract generation we use OpenAI API (GPT-4o) as the main model and instructor to constrain the output [32, 33, 38]
We ran the software on a Lenovo ThinkPad T14 with an 11th Gen Intel® Core™ i5-1145G7 CPU and 16GB of DDR4 RAM. The pipeline took close to two hours to process the WARC files and another two hours to generate the synthetic data from API calls. We expect faster performance in WARC file processing with a better CPU. In addition, the API call pipelines were called sequentially in order to not breach OpenAI API rate limits. With higher rate limits, it is possible to parallelise the calls for better performance. The WARC files are warc.gz files and are taken from Web Archive Singapore. This provided us with a DataFrame containing synthetic titles and abstracts.
We identified six different combinations of prompts and heuristics as stated above. We then used a ranked aggregation scoring method, incorporating Levenshtein distance for the title and BERTScore (using bert-based-cased embeddings), with criteria of maximum BERTScore median, minimum Levenshtein median, and minimum standard deviation for both, to shortlist two specific combinations of prompts and heuristics. Finally, a team of eight trained cataloguers performed manual grading.